# Diff-IP2D: Diffusion-Based Hand-Object Interaction Prediction on Egocentric Videos

### v3

@@ III-C Training
Training objective. The loss function to train the networks in Diff-IP2D contains four parts, including diffusion-related losses, trajectory loss, affordance loss, and an additional regularization term (see Fig. 2). Taking the right side as an example, we use the variational lower bound $\mathcal{L}_{\text{VLB}}^{\scriptstyle{\text{R}}}$ as the diffusion-related losses:
$\textstyle\mathcal{L}_{\text{VLB}}^{\scriptstyle{\text{R}}}=\sum_{s=2}^{S}||\mathbf{z}_{0}^{\scriptstyle{\text{R}}}-f_{\scriptscriptstyle\text{MADT}}(\mathbf{z}_{s}^{\scriptstyle{\text{R}}},s,M_{\text{seq}})||^{2}+||F_{\text{seq}}^{\scriptstyle{\text{R}}}-\hat{F}_{\text{seq}}^{\scriptstyle{\text{R}}}||^{2},$ (1)
where $O_{n}\!=\!f_{\scriptscriptstyle\text{OAH}}(\hat{F}_{\text{seq}}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{R}}})$ is the predicted contact points, and $\mathcal{L}_{\text{KL}}=\frac{1}{2}(-\log\sigma^{2}_{\scriptscriptstyle\text{OAH}}(\hat{F}_{\text{seq}}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{R}}})+\mu^{2}_{\scriptscriptstyle\text{OAH}}(\hat{F}_{\text{seq}}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{R}}})+\sigma^{2}_{\scriptscriptstyle\text{OAH}}(\hat{F}_{\text{seq}}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{R}}})-1)$ is the KL-Divergence regularization for C-VAE, which is scaled by $c=1\text{e}$-$3$. The latent features and predicted hand waypoints are fused by MLP suggested by the previous work [12]. We consider both reconstructed future HOI features $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}[N_{\text{p}}\!+\!1\!:\!N_{\text{p}}\!+\!N_{\text{f}}]$ and anchored past counterparts $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}[0\!:\!N_{\text{p}}]$ compared to [12] as mentioned before. We also notice that the latent feature spaces before and after the denoising diffusion process represent the same “profile” of the input HOI sequence. Therefore, we propose an additional regularization term implicitly linking $F^{\scriptstyle{\text{R}}}_{\text{seq}}$ and $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}$ by hand trajectory prediction:
where $\tilde{H}_{t}^{\scriptstyle{\text{R}}}=f_{\scriptscriptstyle\text{HTH}}(F_{\text{seq}}^{\scriptstyle{\text{R}}}[N_{\text{p}}\!+\!1\!:\!N_{\text{p}}\!+\!N_{\text{f}}])$. Although Eq. (4) does not explicitly contain the term $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}$, the training direction is the same with Eq. (2), thus maintaining training stability. The regularization helps distill HOI state knowledge by building a closer gradient connection constraining the two latent spaces alongside the diffusion process for better optimization. Here we do not use object affordance prediction for regularization because we empirically found that incorporating OAH mitigates training efficiency while the positive effect is not obvious. Finally, we get the total loss $\mathcal{L}_{\text{total}}$, the weighted sum of all the above-mentioned losses to train our proposed Diff-IP2D. Besides, we leverage the importance sampling technique proposed in improved DDPM [58], which promotes the training process focusing more on the steps with relatively large $\mathcal{L}_{\text{total}}$.
Figure 4: Comparison of AR and our iter-NAR prediction.
III-D Inference
In the inference stage, we first sample $F^{\scriptstyle{\text{R}}}_{\text{noise}}$, $F^{\scriptstyle{\text{L}}}_{\text{noise}}\in\mathbb{R}^{N_{\text{f}}\times a}$ from a standard Gaussian distribution, which is then concatenated with $F^{\scriptstyle{\text{R}}}_{\text{seq}}$, $F^{\scriptstyle{\text{L}}}_{\text{seq}}\in\mathbb{R}^{N_{\text{p}}\times a}$ along the time axis to generate $\mathbf{z}_{S}^{\scriptstyle{\text{R}}}$ and $\mathbf{z}_{S}^{\scriptstyle{\text{L}}}$. Then we use MADT to predict $\mathbf{z}_{0}^{\scriptstyle{\text{R}}}$ and $\mathbf{z}_{0}^{\scriptstyle{\text{L}}}$ based on DDIM sampling [59]. Note that we anchor the past part of reparameterized $\mathbf{z}_{s}$ as the fixed condition in every step of the inference process following Gong et al. [26]. Finally, the generated $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}$ and $\hat{F}^{\scriptstyle{\text{L}}}_{\text{seq}}$ are used to predict future hand waypoints and contact points by $f_{\scriptscriptstyle\text{HTH}}(\cdot)$ and $f_{\scriptscriptstyle\text{OAH}}(\cdot)$ as mentioned before. It can be seen from the inference stage that Diff-IP2D can be regarded as an iter-NAR model in the latent feature space.
Takeaways. Compared to the state-of-the-art baselines in an autoregressive manner, Diff-IP2D shifts the limited iterations along the time axis to sufficient iterations in the diffusion denoising direction, as shown in Fig. 4. This alleviates the accumulated artifacts caused by the limited iterations in the time dimension, and maintains bidirectional constraints (inherent in MADT) among the sequential features to generate future HOI states in parallel. Bidirectional constraints respect the fact that possible final HOI states also affect prior hand trajectories beyond temporal causality. We argue that this new paradigm provides a deeper understanding of high-level human intention for more accurate HOI prediction.
IV Experiments
IV-A Experimental setups
Datasets. Following the previous work [12], we utilize three public datasets, Epic-Kitchens-55 (EK55) [60], Epic-Kitchens-100 (EK100) [61], and EGTEA Gaze+ (EG) [11]. For the EK55 and EK100 datasets, we sample past $N_{\text{p}}=10$ frames (2.5 s) to forecast HOI states in future $N_{\text{f}}=4$ frames (1.0 s), both at 4 FPS. As to the EG dataset, $N_{\text{p}}=9$ frames (1.5 s) are used for $N_{\text{f}}=3$ HOI predictions (0.5 s) at 6 FPS. All the training and test splits are obtained following [12].
Diff-IP2D configuration. MFE extracts the hand, object, and global feature vectors all with the size of $512$ for each input image. For the EK55 and EK100 datasets, the outputs of SOFM $F^{\scriptstyle{\text{R}}}_{\text{seq}}$, $F^{\scriptstyle{\text{L}}}_{\text{seq}}$ have the size of ${14\times 512}$ for training and ${10\times 512}$ for inference. For the EG dataset, $F^{\scriptstyle{\text{R}}}_{\text{seq}}$, $F^{\scriptstyle{\text{L}}}_{\text{seq}}$ are ${9\times 512}$ for training and ${12\times 512}$ for inference. As to the diffusion process, the total number of steps $S$ is set to $1000$. The square-root noise schedule in Diffusion-LM [62] is adopted here for the forward diffusion process. MADT has $6$ Transformer layers (Fig. 3) for denoising, where the embedding dimension is $512$, the number of heads is set to $4$, and the intermediate dimension of the feed-forward layer is set to $2048$. Motion Encoder linearly projects each homography matrix to an egomotion feature vector of $512$. We use an MLP with hidden dimensions $256$ and $64$ to predict the hand waypoints as HTH, and a C-VAE containing an MLP with a hidden dimension $512$ to predict contact points as OAH. For training Diff-IP2D, we use AdamW optimizer [63] with a learning rate 2e-4. All the modules are trained for 30 epochs with a batch size of 8 on 2 A100 GPUs. In the reference stage, we generate the $10$ candidate samples for each prediction.
Evaluation metrics. Following the previous works [14, 12, 13], we use Final Displacement Error (FDE) to evaluate hand trajectory prediction performance. Considering the general knowledge of “post-contact trajectories” extracted from human videos is potentially beneficial to robot manipulation [1, 18], we additionally extend the metric Average Displacement Error to Weighted Displacement Error (WDE):
where $D(\cdot)$ denotes the L2 distance function and later waypoints contribute to larger errors. We select the mean error among the 10 samples for each trajectory prediction. As to object affordance prediction, we use Similarity Metric (SIM) [64], AUC-Judd (AUC-J) [65], and Normalized Scanpath Saliency (NSS) [66] as evaluation metrics. We use all 10 contact point candidates to compute them.
Moreover, we propose a novel object-centric protocol to jointly evaluate the two prediction tasks. We first calculate the averaged hand waypoints $\bar{H}^{\scriptstyle{\text{R}}}_{t}$ and $\bar{H}^{\scriptstyle{\text{L}}}_{t}$ for each future timestamp from multiple samples. Then we select the waypoint closest to each predicted contact prediction $O_{n}$ as an additional “interaction point”, which can be formulated by:
$\textstyle\bar{H}_{n}^{\text{ip}}=\textrm{min}_{{\scriptstyle{\text{R}}},{\scriptstyle{\text{L}}},t}D(\bar{H}_{t},O_{n}).$ (6)
Ultimately, the joint hotspot is predicted using $\{\bar{H}_{n}^{\text{ip}}\cup O_{n}\}_{n=1}^{N_{\text{o}}}$. This protocol comprehensively considers object-centric attention since HOI changes object states and hand waypoints must have a strong correlation with object positions. Note that we also use the quantitative metrics same as the ones for object affordance prediction, which are denoted as SIM, AUC-J, and NSS.
Figure 5: Visualization of hand trajectory prediction on Epic-Kitchens. The waypoints from ground-truth, Diff-IP2D, and the second-best baseline [12] are connected by red, white, and blue dashed lines.

### v4

@@ III-C Training
Training objectives. The loss function training Diff-IP2D comprises four components: diffusion-related losses, trajectory loss, affordance loss, and a regularization term (see Fig. II). Taking the right side as an example, we use the variational lower bound $\mathcal{L}_{\text{VLB}}^{\scriptstyle{\text{R}}}$ as diffusion-related losses:
$\textstyle\mathcal{L}_{\text{VLB}}^{\scriptstyle{\text{R}}}=\sum_{s=2}^{S}||\mathbf{z}_{0}^{\scriptstyle{\text{R}}}-f_{\scriptscriptstyle\text{MADT}}(\mathbf{z}_{s}^{\scriptstyle{\text{R}}},s,M_{\text{seq}})||^{2}$
$+||F_{\text{seq}}^{\scriptstyle{\text{R}}}-\hat{F}_{\text{seq}}^{\scriptstyle{\text{R}}}||^{2},$ (1)
where $O_{n}\!=\!f_{\scriptscriptstyle\text{OAH}}(\hat{F}_{\text{seq}}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{R}}})$ is the predicted contact points, and $\mathcal{L}_{\text{KL}}=\frac{1}{2}(-\log\sigma^{2}_{\scriptscriptstyle\text{OAH}}(\hat{F}_{\text{seq}}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{R}}})+\mu^{2}_{\scriptscriptstyle\text{OAH}}(\hat{F}_{\text{seq}}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{R}}})+\sigma^{2}_{\scriptscriptstyle\text{OAH}}(\hat{F}_{\text{seq}}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{R}}})-1)$ is the KL-Divergence regularization for C-VAE, which is scaled by $c=1e\text{-}3$. The latent features and predicted hand waypoints are fused by MLP suggested by the previous work [10]. We consider both reconstructed future HOI features $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}[N_{\text{p}}\!+\!1\!:\!N_{\text{p}}\!+\!N_{\text{f}}]$ and anchored past counterparts $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}[0\!:\!N_{\text{p}}]$ compared to [10] as mentioned before. We also notice that the latent feature spaces before and after the denoising diffusion process represent the same “profile” of the input HOI sequence. Therefore, we propose an additional regularization term implicitly linking $F^{\scriptstyle{\text{R}}}_{\text{seq}}$ and $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}$ by hand trajectory prediction:
where $\tilde{H}_{t}^{\scriptstyle{\text{R}}}=f_{\scriptscriptstyle\text{HTH}}(F_{\text{seq}}^{\scriptstyle{\text{R}}}[N_{\text{p}}\!+\!1\!:\!N_{\text{p}}\!+\!N_{\text{f}}])$. Although Eq. (4) does not explicitly contain the term $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}$, the training direction is the same with Eq. (2), thus maintaining training stability. The regularization helps distill HOI state knowledge by building a closer gradient connection constraining the two latent spaces alongside the diffusion process for better optimization. Here we do not use object affordance prediction for regularization because we empirically found that incorporating OAH mitigates training efficiency while the positive effect is not obvious. Sec. 3 in the supplementary material provides more detailed clarification about the motivation of our proposed regularization strategy. Finally, we get the total loss $\mathcal{L}_{\text{total}}$, the weighted sum of all the above-mentioned losses to train our proposed Diff-IP2D. Besides, we leverage the importance sampling technique proposed in improved DDPM [54], which promotes the training process focusing more on the steps with relatively large $\mathcal{L}_{\text{total}}$.
Figure IV: Comparison of AR prediction with forward-only constraints and our iter-NAR prediction with both forward and backward constraints. The yellow star represents the starting point of trajectory prediction, and the orange triangle represents the interaction point with the target object.
III-D Inference
Prediction pipeline. In the inference stage, we first sample $F^{\scriptstyle{\text{R}}}_{\text{noise}}$, $F^{\scriptstyle{\text{L}}}_{\text{noise}}\in\mathbb{R}^{N_{\text{f}}\times a}$ from a standard Gaussian distribution, which are then concatenated with $F^{\scriptstyle{\text{R}}}_{\text{seq}}$, $F^{\scriptstyle{\text{L}}}_{\text{seq}}\in\mathbb{R}^{N_{\text{p}}\times a}$ along the time axis to generate $\mathbf{z}_{S}^{\scriptstyle{\text{R}}}$ and $\mathbf{z}_{S}^{\scriptstyle{\text{L}}}$. Then we use MADT to predict $\mathbf{z}_{0}^{\scriptstyle{\text{R}}}$ and $\mathbf{z}_{0}^{\scriptstyle{\text{L}}}$ based on DDIM sampling [55]. Note that we anchor the past part of reparameterized $\mathbf{z}_{s}$ as the fixed condition in every step of the inference process following Gong et al. [24]. Finally, the generated $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}$ and $\hat{F}^{\scriptstyle{\text{L}}}_{\text{seq}}$ are used to predict future hand waypoints and contact points by $f_{\scriptscriptstyle\text{HTH}}(\cdot)$ and $f_{\scriptscriptstyle\text{OAH}}(\cdot)$ as mentioned before. It can be seen from the inference stage depicted in Fig. IV that Diff-IP2D is iter-NAR with bidirectional constraints in the latent feature space (also refer to Sec. 2 of the supplementary material).
Temporal enhancement. We propose an optional temporal enhancement strategy to further improve the prediction performance of Diff-IP2D inspired by [56]. As Fig. V illustrates, we can incorporate an additional autoregressive setup at the end of our proposed iter-NAR paradigm. This helps to improve the smoothness of predicted hand trajectories and simultaneously refine affordance prediction. Specifically, we exploit a sliding window to predict hand waypoints and contact points at each future timestamp. Then an exponential weighting scheme $w_{t}=\text{exp}(-u*t)$ is adopted to aggregate these results, where $u$ is constant. Our proposed temporal enhancement does not affect the concrete pattern of each prediction (e.g. the fixed number of input frames), and thus there is no need to integrate this step-wise operation in the training stage. We denote Diff-IP2D with temporal enhancement as Diff-IP2D, which is suited to applications with low real-time requirements since multiple denoising operations should be performed to obtain ultimate prediction results. Recall that Diff-IP2D follows the iter-NAR paradigm considering bidirectional constraints in each HOI process, and Diff-IP2D still keeps bidirectional constraints during denoising operation at each timestamp. The difference is that Diff-IP2D enhances forward constraints to improve trajectory smoothness and reduce affordance uncertainty by weighting step-wise denoised HOI.
Figure V: Temporal enhancement aggregates multiple prediction results to enhance forward constraints while still keeping backward constraints within each denoising process.
@@ III-D Inference
IV Experiments
IV-A Experimental Setups
Datasets. Following the previous work [10, 11], we utilize four public datasets, Epic-Kitchens-55 (EK55) [59], Epic-Kitchens-100 (EK100) [60], EGTEA Gaze+ (EG) [9], and EgoPAT3D-DT [61, 11]. For the EK55 and EK100 datasets, we sample past $N_{\text{p}}=10$ frames (2.5 s) to forecast HOI states in future $N_{\text{f}}=4$ frames (1.0 s), both at 4 FPS. As to the EG dataset, $N_{\text{p}}=9$ frames (1.5 s) are used for $N_{\text{f}}=3$ HOI predictions (0.5 s) at 6 FPS. All the training and test splits of EK55, EK100, and EG are obtained following [10]. More details can be found in Sec. 4 of the supplementary material. The experiments conducted on EgoPAT3D-DT are also provided in Sec. 5-D of the supplementary material.
Diff-IP2D configuration. MFE extracts the hand, object, and global feature vectors all with the size of $512$ for each input image. For the EK55 and EK100 datasets, the outputs of SOFM $F^{\scriptstyle{\text{R}}}_{\text{seq}}$, $F^{\scriptstyle{\text{L}}}_{\text{seq}}$ have the size of ${14\times 512}$ for training and ${10\times 512}$ for inference. For the EG dataset, $F^{\scriptstyle{\text{R}}}_{\text{seq}}$, $F^{\scriptstyle{\text{L}}}_{\text{seq}}$ are ${9\times 512}$ for training and ${12\times 512}$ for inference. As to the diffusion process, the total number of steps $S$ is set to $1000$. The square-root noise schedule in Diffusion-LM [62] is adopted here for the forward diffusion process. MADT has $6$ Transformer layers (Fig. III) for denoising, where the embedding dimension is $512$, the number of heads is set to $4$, and the intermediate dimension of the feed-forward layer is set to $2048$. Motion Encoder linearly projects each homography matrix to an egomotion feature vector of $512$. We use an MLP with hidden dimensions $256$ and $64$ to predict the hand waypoints as HTH, and a C-VAE containing an MLP with a hidden dimension $512$ to predict contact points as OAH. For training Diff-IP2D, we use AdamW optimizer [63] with a learning rate 2e-4. All the modules are trained for 30 epochs with a batch size of 8 on 2 A100 GPUs. In the reference stage, we generate $10$ candidate samples for each prediction.
Evaluation metrics. Following previous works [12, 10, 11], we use Final Displacement Error (FDE) to evaluate hand trajectory prediction performance. Considering the general knowledge of “post-contact trajectories” extracted from human videos is potentially beneficial to robot manipulation [1, 16], we additionally extend the metric Average Displacement Error to Weighted Displacement Error (WDE):
where $D(\cdot)$ denotes the L2 distance function and later waypoints contribute to larger errors. We select the mean error among the 10 candidate samples for each trajectory prediction. As to object affordance prediction, we use Similarity Metric (SIM) [64], AUC-Judd (AUC-J) [65], and Normalized Scanpath Saliency (NSS) [66] as evaluation metrics. We use all 10 contact point candidates to compute them.
Moreover, we exploit an object-centric protocol to jointly evaluate the two prediction tasks. We first calculate the averaged hand waypoints $\bar{H}^{\scriptstyle{\text{R}}}_{t}$ and $\bar{H}^{\scriptstyle{\text{L}}}_{t}$ for each future timestamp from multiple samples. Then we select the waypoint closest to each predicted contact prediction $O_{n}$ as an additional possible contact point. The joint hotspot is predicted with the additional contact points and $O_{n}$. This comprehensively considers object-centric attention since HOI changes object states and hand waypoints must have a strong correlation with object positions. Here we use the quantitative metrics same as the ones for affordance prediction, denoted as SIM, AUC-J, and NSS.
Figure VI: Visualization of hand trajectory prediction on Epic-Kitchens. Waypoints from GT, Diff-IP2D, and the second-best baseline [10] are connected by red, white, and blue dashed lines.
Figure VII: Visualization of object affordance prediction on Epic-Kitchens. Contact points from GT, Diff-IP2D, and the SOTA baseline OCT [10] are represented by red, white, and blue dots respectively. For clearer illustration, we additionally put a fixed Gaussian with each contact point as the center.
We first present the evaluation results on hand trajectory prediction. As Tab. I depicts, our proposed Diff-IP2D outperforms all the baselines on both EK55 and EK100 on WDE and FDE. This is mainly achieved by the devised iter-NAR paradigm of Diff-IP2D alleviating degeneration in AR baselines, as well as the egomotion guidance. The visualization of hand prediction results in Fig. VI shows that Diff-IP2D can better capture the camera wearer’s intention (such as putting the food in the bowl) and generate more reasonable future trajectories even if lacking past observations for hands (such as reaching out towards the table). Besides, Diff-IP2D can predict a good final hand position despite a large shift in the early stage (Fig. VI, bottom right), owing to our parallel generation with bidirectional constraints. When directly transferring the models trained on Epic-Kitchens to the unseen EG dataset, Diff-IP2D still improves the second-best baselines by 7.0% and 15.3% on WDE and FDE respectively. This reveals the solid generalization capability of our Diff-IP2D across different environments.
The comparison results of object affordance prediction are also shown in Tab. I. Our proposed Diff-IP2D predicts the hotspots with larger SIM, AUC-J, and NSS compared to all the baselines on both Epic-Kitchens data and unseen EG data. Fig. VII illustrates the predicted contact points with minimum distances to the ground-truth ones. Our proposed method focuses more on objects of interest considering the features of the holistic interaction and potential hand trajectories, and therefore grounds the contact points closer to the ground-truth labels than the counterparts of the baseline.
TABLE III: Ablation study on supervision signals
approach hand trajectory object affordance joint evaluation
WDE $\downarrow$ FDE $\downarrow$ SIM $\uparrow$ AUC-J $\uparrow$ NSS $\uparrow$ ${}^{*\!}\uparrow$ ${}^{*\!}\uparrow$ ${}^{*\!}\uparrow$
Diff-IP2D w/o diff. 0.480 0.201 0.142 0.624 0.406 0.189 0.634 0.764
Diff-IP2D w/o reg. 0.430 0.195 0.205 0.718 0.821 0.180 0.692 0.722
Diff-IP2D 0.407 0.187 0.211 0.736 0.917 0.204 0.727 0.844
TABLE IV: Diff-IP2D vs. Diff-IP2D with temporal enhancement
approach EK55 EK100
Diff-IP2D 0.411 0.181 0.226 0.725 0.980 0.407 0.187 0.211 0.736 0.917
Diff-IP2D 0.388 0.172 0.230 0.735 0.992 0.395 0.179 0.217 0.744 0.930
improvement 5.6% 5.0% 1.8% 1.4% 1.2% 2.9% 4.3% 2.8% 1.1% 1.4%
IV-C Joint Evaluation on HOI Prediction
We further compare Diff-IP2D with the other two joint prediction baselines, FHOI [12] and OCT [10], using the object-centric protocol. The video clips containing both ground-truth hand waypoints and contact points are used for evaluation in this experiment. The results are also shown in Tab. I, which indicates that our proposed Diff-IP2D can generate the best object-centric HOI predictions considering the two tasks concurrently on both Epic-Kitchens and unseen EG data. The results also suggest that Diff-IP2D outperforms the baselines on object-centric HOI prediction by focusing more attention on the target objects and predicting reasonable hand trajectories around them.
TABLE V: Ablation study on the denoising model
denoising EK55 EK100
WDE $\downarrow$ FDE $\downarrow$ WDE $\downarrow$ FDE $\downarrow$
separate models 0.425 0.189 0.412 0.189
unified model 0.411 0.181 0.407 0.187
IV-D Ablation Studies
Egomotion guidance. We first ablate the egomotion features used to guide MADT denoising on the EK55 and EK100 datasets. Here we replace the MHCA in MADT with a multi-head self-attention module (MHSA) to remove the egomotion guidance while keeping the same parameter number. The experimental results in Tab. II show that the guidance of motion features improves our proposed diffusion-based paradigm noticeably on both hand trajectory prediction and object affordance prediction. This is achieved by narrowing the two gaps caused by 2D-3D ill-posed problem and view difference mentioned in Sec. I. Note that the egomotion guidance is more significant on the EK100 dataset than on the EK55 dataset. The reason could be that EK100 has a larger volume of training data incorporating more diverse egomotion patterns than EK55, leading to a model that can capture human dynamics better.
Supervision signals. We provide an additional ablation study on diffusion-related losses, and the regularization term which links $\{F^{\scriptstyle{\text{R}}}_{\text{seq}},F^{\scriptstyle{\text{L}}}_{\text{seq}}\}$ and $\{\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}},\hat{F}^{\scriptstyle{\text{L}}}_{\text{seq}}\}$. The experimental results on the EK100 dataset are shown in Tab. III. Diff-IP2D without diffusion-related losses witness significant performance degradation due to a lack of dense supervision signals in the latent space beyond sparse constraints in the image plane. We thus argue that high-level dense supervision aligns the model with high-level human intentions for better prediction performance. Tab. III also shows that our regularization strategy remarkably enhances prediction performance on both hand trajectories and object affordances even if it is only used to link the latent space with hand trajectory prediction. More analysis of the regularization can be found in Sec. 3 of the supplementary material.
Denoising model. We conduct a baseline by instantiating two separate MADT models for right and left sides, and denoise side-oriented HOI features of two sides respectively. Tab. V shows that denoising with a unified MADT for both sides outperforms separate denoising. This suggests that side-oriented denoising with a unified model rather than separate ones helps to capture potential correlations between right and left sides, leading to higher prediction accuracy.
Temporal enhancement. We ablate our proposed temporal enhancement by comparing our vanilla Diff-IP2D to Diff-IP2D mentioned in Sec. III-D on EK55 and EK100. Tab. IV shows that the proposed temporal enhancement improves Diff-IP2D prediction accuracy. Note that temporal enhancement is an optional strategy since Diff-IP2D has already outperformed baselines. We advocate using Diff-IP2D when there is no requirement of high efficiency for better HOI prediction performance.
IV-E Key Findings
Iter-NAR paradigm. Compared to the SOTA baselines in an autoregressive manner shown in Tab. I, Diff-IP2D shifts limited iterations along the time axis to sufficient iterations in the diffusion denoising direction, as Fig. IV shows. This alleviates accumulated artifacts caused by the limited iterations with forward-only constraints in the time dimension, and maintains bidirectional constraints (inherent in MADT) among sequential features to generate future HOI states in parallel. Bidirectional constraints respect spatial causality where possible final HOI states also affect prior hand trajectories, without losing temporal causality. We argue that this new paradigm provides a deeper understanding of high-level human intention for more accurate HOI prediction.
Concurrent motion capture. Diff-IP2D with egomotion guidance is inherently suited to egocentric views because it concurrently captures hand/object movements and the camera wearer’s egomotion patterns (homography) by the proposed MADT. It respects the fact that the changes of hand/object locations within the field of vision are entangled with human head motion following specific intentions in different activities. Tab. II shows how egomotion affects HOI prediction positively. Additionally, our paradigm may also be beneficial to HOI prediction in non-egocentric views. For instance, if the human body egomotion can be captured in a third-person perspective, our model can also associate it with hand movements, thereby achieving better HOI prediction. We will explore this topic in future work.
Side-oriented denoising with a unified model. Diff-IP2D uses SOFM to separately fuse and denoise features for the left and right sides. We argue that the optimization directions for left-hand and right-hand HOI are different considering their different motion patterns and approaches to active objects. Therefore, separate noise sampling should be applied for the respective side in the training process. Notably, as Tab. V presents, we advocate using a unified model MADT to denoise for both sides with different sampled noises since we encourage it to capture the potential correlation between sides during one interaction process.
Dense supervision signals. We extend sparse supervision signals from explicit ground-truth hand trajectories and contact points to the latent space. Specifically, we incorporate reconstructed implicit HOI features into training losses in our diffusion-based scheme, alongside explicit ground-truth supervision. We argue that this high-level dense supervision aligns the model with high-level human intentions, leading to better HOI prediction as Tab. III shows.
V Conclusion
In this paper, we propose a novel hand-object interaction prediction method Diff-IP2D. It implements the devised denoising diffusion in the latent space under our proposed egomotion guidance, and jointly predicts future hand trajectories and object affordances with recovered latents on 2D egocentric videos. Experimental results validate that Diff-IP2D dominates the existing baselines on extended metrics, suggesting promising applications in artificial intelligence systems. We hope the takeaways about the iter-NAR paradigm, concurrent motion capture, side-oriented denoising with a unified model, and dense supervision signals in Diff-IP2D could inspire future HOI prediction research.
Supplementary Material
1 Motion-Related Gaps and Egomotion Homography
In this section, we provide a detailed analysis for filling the motion-related gaps in hand-object interaction (HOI) prediction mentioned in Sec. I of the main text with the egomotion homography. To narrow the view gap between the last observation and the other observations, homography works as a bridge to connect the pixel positions $\mathbf{p}_{0}$, $\mathbf{p}_{t}\in\mathbb{R}^{2}$ of one 3D hand waypoint/contact point on $I_{t}$ ($t\leq 0$) and $I_{0}$, which can be represented by $\mathbf{p}_{0}=M_{t}\mathbf{p}_{t}$. We let the denoising network be aware of the egomotion features $E_{t}$ encoded from $M_{t}$ and enable it to capture the above-mentioned transformation when predicting future hand trajectories and contact points on the last observed image as a canvas.
For the 2D-3D gaps, we first discover the relationship between 2D pixel movements and 3D hand movements. For a 3D point that moves from $\mathbf{P}_{t}\in\mathbb{R}^{3}$ in the camera coordinate system at timestamp $t$ ($t\leq 0$) to $\mathbf{P}_{0}\in\mathbb{R}^{3}$ in the camera coordinate system at timestamp $t=0$, we first project them to the image plane by $\mathbf{p}_{t}=K\mathbf{P}_{t}$ and $\mathbf{p}_{0}=K\mathbf{P}_{0}$, where $K$ is the intrinsic parameters. Then we transform $\mathbf{p}_{t}$ to the last canvas image by $\mathbf{p}^{\prime}_{t}=M_{t}\mathbf{p}_{t}$. The 2D pixel movement on the last image can be formulated as:
$\mathbf{p}_{0}-\mathbf{p}^{\prime}_{t}=K\mathbf{P}_{0}-M_{t}\mathbf{p}_{t}=K\mathbf{P}_{0}-M_{t}K\mathbf{P}_{t}.$
Therefore, the 3D action ($\mathbf{P}_{t}\rightarrow\mathbf{P}_{0}$) uniquely corresponds to the 2D pixel movement ($\mathbf{p}_{t}\rightarrow\mathbf{p}_{0}$) once $K$ and $M_{t}$ are both determined. Since $K$ is a constant for each video clip, only $M_{t}$ changing along the time axis determines the spatial relationship between observations. Therefore, we enable our proposed model aware of egomotion by encoding $M_{t}$ to a feature vector absorbed by multi-head cross attention of Motion-Aware Denoising Transformer as mentioned in Sec. III-B of the main text, narrowing the gap between 2D pixel movement and 3D actions. Note that we do not utilize SE(3) here due to scale-agnostic estimation with only 2D images as input.
2 Iterative Non-Autoregressive Paradigm vs. Autoregressive Paradigm
Our proposed Diff-IP2D is an iterative non-autoregressive (iter-NAR) model, showing better HOI prediction performance compared to the state-of-the-art methods [10, 11] with the autoregressive (AR) paradigm. AR models reason about the next HOI state only according to the previous steps (Fig. I(a) in the main text), leading to the forward-only constraint. They overlook the backward constraint which we think is also important for HOI prediction. We provide an example in Fig. 2 to further explain the significance of the backward constraint. The human hand generally picks up a cup (Fig. 2(a)) by its handle because side-gripping by the handle is more stable and allows for a faster target approach than other ways. It is more likely for a hand to approach the cup from the side (red arrow in Fig. 2(b)) than from the top (green arrow in Fig. 2(b)) in the near future. Consequently, the final state of the future HOI can be approximately determined, which dictates the hand movement toward the cup, thereby establishing potential backward constraints on spatial causality. Therefore, we argue that HOI prediction should be modeled as the non-autoregressive process considering the bidirectional constraints within the holistic sequence, rather than the autoregressive process with only forward constraints on temporal causality.
Figure 1: Autoregressive generation vs. parallel generation.
Figure 2: An example to clarify our motivation to propose an iter-NAR paradigm considering bidirectional constraints. The hand waypoints from ground-truth labels and HOI prediction approaches are connected by blue and white dashed lines respectively. Note that we reverse the RGB values of each image to display the arm’s positions more clearly. There is a lack of backward constraints in AR-based USST [11], leading to a shorter predicted trajectory (almost curled up into a point) and larger accumulated displacement errors. In contrast, our Diff-IP2D with iter-NAR paradigm is potentially guided by final HOI states, and thus predicts more accurate hand trajectories following both spatial causality and temporal causality.
TABLE 1: Joint evaluation results in the ablation study on egomotion guidance
      approach       EK55       EK100
      ${}^{*\!}\uparrow$       ${}^{*\!}\uparrow$       ${}^{*\!}\uparrow$       ${}^{*\!}\uparrow$       ${}^{*\!}\uparrow$       ${}^{*\!}\uparrow$
      Diff-IP2D*       0.216       0.718       0.842       0.198       0.712       0.778
      Diff-IP2D       0.222       0.730       0.888       0.204       0.727       0.844
      improvement       2.8%       1.7%       5.5%       3.0%       2.1%       8.5%